You are here: TWiki > SoftwareTeam Web>Meetings>2011-09-17 (28 Sep 2011, JamesWeichel)

OSG Software Team Meeeting Minutes, 19-Sep-2011

Attending

  • Cartwright, Tim
  • Dost, Jeff
  • Kronenfeld, Scot
  • Mambello, Marco
  • Roy, Alain
  • Sim, Alex
  • Strain, Doug
  • Thapa, Suchandra

Apologies

  • Levshina, Tanya
  • Sfiligoi, Igor
  • Selmeci, Mat
  • Sharma, Neha

Globus 5.2

Globus 5.2 is fully ready for testing. The last major blocker was a problem with init scripts, and Mat has fixed that.

Signed RPMs

All RPMs are now signed. You should notice no different--install as usual.

CA Certificates

no-ca-certs has been replaced by empty-ca-certs to work properly with yum's crazy dependency resolution mechanism.

State of testing

Mat: Mat did an initial test of the osg-ce and found the init problems. He'll do further testing.

Tim: Tests caught minor problems that have been addressed.

Scot: Our Condor still doesn't work. Alain is working with Globus folks to address the problem. [Update from 20-Sep: problem has been found, deciding on a fix.]

Doug: Haven't tried the new xrootd-dsi, was working on cemon. Will do short smoke test to ensure that xrootd-dsi and GridFTP? still work.

Marco: Haven't tried it yet, will try today.

Suchandra: Did worker node client on Friday: got it installed but from osg-testing. Will try again with osg-development.

Status reports

  • Alain: Various management, empty-ca-certs.
  • Alex: Bestman Assessmnet last week. Still working with Bestman problem with OU's problem with NFS. Working with OpenJDK? 1.7 update & dependent libraries. Hoping to release new version with new dependencies by the end of hte month.
  • Doug: Xrootd meeting and Bestman Assessment. Cemon & dependent packages built last week, but haven't tested it yet.
  • Jeff: Last week got sidetracked with glidein stuff because Igor was gone and there was a serious bug. Will make up lost time from last week to work with HDFS migration.
  • Marco: Will work on testing this week & CA certificates probe. Last week fixed osg-site-web-page
  • Mat: finished up GT 5.2 update & rebuilds of software against Globus 5.2. Also added signed packages.
  • Scot: RSV works with Condor-G as long as it's the one from the Condor Team, not the one from our repo. Working on fixing globus-job-run support. Ready for testing. Doug asked about a condor cron bug: it's been fixed. Also fixed Apache in Pacman. Added osg-cert-scripts to the VDT's repo.
  • Suchandra: XRootd meeting & Bestman assessment last week. Testing worker node. Will help out testing the CE this week. Will configure CEMon soon. Will be on vacation next week.
  • Tim: Rolled out test stuff last week. Worked on Debian 6 work for LIGO as well.
Topic revision: r3 - 28 Sep 2011 - 20:07:15 - JamesWeichel
Hello, TWikiGuest
Register

 
TWIKI.NET

TWiki | Report Bugs | Privacy Policy

This site is powered by the TWiki collaboration platformCopyright by the contributing authors. All material on this collaboration platform is the property of the contributing authors..